# Zhihui Zhu

### v1

@@ C.1 Proof of Proposition 3
where the second equality follows from $m_{k}\alpha_{k}/m=\alpha$ according to (11). This implies $\bm{0}$ is a local maximizer of $f_{k}(\bm{Z}_{k})$. Suppose to the contrary that $r_{k}>0$ for all $k\in[K]$. For each $\bm{Z}\in\cal Z$, using Lemma 1 with $\bm{Z}_{k}^{T}\bm{Z}_{l}=\bm{0}$ for all $k\neq l$, (14), and (31), we have
@@
C.2 Proof of Proposition 4
Proof of Proposition 4.
@@ C.2 Proof of Proposition 4
According to Lemma 5, we can see that $\bm{Z}$ is a critical point with curvature if and only if $\bm{Z}\bm{O}$ is a critical point with the same curvature for each $\bm{O}=\mathrm{BlkDiag}\left(\bm{O}_{1},\dots,\bm{O}_{K}\right)$ with $\bm{O}_{k}\in\mathcal{O}^{m_{k}}$ for all $k\in[K]$. According to the SVD of $\bm{Z}_{k}$ in (33), we can take $\bm{O}_{k}=\bm{Q}_{k}$ for each $k\in[K]$. Therefore, it suffices to study $\bm{Z}_{k}=\bm{P}_{k}\bm{\Sigma}_{k}$ for each $k\in[K]$. Substituting this into (24) in Definition 1 gives
@@
This, together with (2), Lemma 4, and $\bm{Z}^{(i)^{T}}\bm{Z}^{(j)}=\bm{0}$, yields that
@@
$\quad-\alpha\alpha_{k}\left(1-\frac{\alpha_{k}}{\alpha_{k}\beta_{i}^{2}+1}\|\bm{Z}_{k}^{(i)^{T}}\bm{u}\|^{2}\right)\frac{\beta_{i}^{2}\|\bm{q}_{k}\|^{2}}{\alpha_{k}\beta_{i}^{2}+1},$
@@
$\quad\sum_{k=1}^{K}\nabla^{2}R_{c}\left(\bm{Z}_{k}^{(i)}\right)\left[\bm{D}_{k}^{(i)},\bm{D}_{k}^{(i)}\right]$
@@
$\quad-\frac{\alpha\alpha_{k}}{(\alpha_{k}\beta_{i}^{2}+1)^{2}}\sum_{k=1}^{K}\left(\bm{u}^{T}\bm{Z}_{k}^{(i)}\bm{q}_{k}\right)^{2}+\frac{\alpha\alpha_{k}^{2}\beta_{i}^{2}}{\left(\alpha_{k}\beta_{i}^{2}+1\right)^{2}}\sum_{k=1}^{K}\|\bm{q}_{k}\|^{2}\|\bm{Z}_{k}^{(i)^{T}}\bm{u}\|^{2}$
@@
Appendix D Proofs in Section 3
D.1 Proof of Theorem 1
@@
(i) Suppose that each block of $\bm{Z}$ satisfying (7), (a), (b), and (c). It directly follows from (i) in Proposition 3 that $\bm{Z}$ is a local maximizer. Conversely, suppose that $\bm{Z}$ is a local maximizer. According to (17), (18), (19), (ii) in Proposition 3 and Proposition 4, if $\bm{Z}\in\mathcal{Z}^{c}\cup\mathcal{Z}_{2}$, then $\bm{Z}$ is a strict saddle point. This, together with $\mathcal{X}=\mathcal{Z}^{c}\cup\mathcal{Z}_{1}\cup\mathcal{Z}_{2}$ and the fact that $\bm{Z}$ is a local maximizer, implies that $\bm{Z}\in\mathcal{Z}_{1}$. Using this, (19), and Proposition 2 yields that $\bm{Z}$ satisfying (7), (a), (b), and (c).
(ii) According to (i) in Theorem 1, suppose that the $k$-th block of a local maximizer $\bm{Z}$ admits the decomposition in (7) satisfying (a), (b), and (c) for all $k\in[K]$. This, together with (C.1) in the proof of Proposition 3, yields that
@@
D.2 Proof of Proposition 1
To prove Proposition 1, we first need to characterize the global optimal solution set of Problem (4).
@@ D.2 Proof of Proposition 1
According to Lemma 4, we have
@@
$\max_{\sigma_{k,1},\dots,\sigma_{k,r_{k}}}\sum_{i=1}^{r_{k}}\log\left(1+\alpha\sigma_{k,i}^{2}\right)-\sum_{i=1}^{r_{k}}\frac{m_{k}}{m}\log\left(1+\alpha_{k}\sigma_{k,i}^{2}\right)\qquad\mathrm{s.t.}\ \sum_{i=1}^{r_{k}}\sigma_{k,i}^{2}=m_{k}.$
@@
$\max_{x_{1},\dots,x_{r_{k}}}g(\bm{x})$ $:=\sum_{i=1}^{r_{k}}\log\left(1+\alpha x_{i}\right)-\sum_{i=1}^{r_{k}}\frac{1}{K}\log\left(1+\alpha_{k}x_{i}\right)\qquad$
@@
This, together with Lemma 6 and (11), yields that the optimal solution for each $k\in[K]$ is
@@
(i) Suppose that $m\leq d$. Then, we have $r_{k}\leq m/K$ for each $k\in[K]$ and $\sum_{k=1}^{K}r_{k}\leq m$. This, together with (65) and Lemma 7, implies that $r_{k}=m/K$, and thus $x_{i}^{*}=1$ for all $i\in[r_{k}]$ and $k\in[k]$.
@@
Based on the above proposition, we are ready to prove Proposition 1.
Proof of Proposition 1.
Let $\hat{\bm{Z}}=[\bm{Z}_{1},\dots,\bm{Z}_{k}]$ denote the optimal solution of Problem (4). According to Proposition 5, it suffices to study the following two cases.
(i) Suppose that $m<d$. Using this and (61), we have $\hat{\bm{Z}}_{k}=\bm{U}_{k}\bm{V}_{k}^{T}$ and $\hat{r}_{k}=m/K$ for each $k\in[K]$. Moreover, according to Theorem 1, if $m_{k}=m/K$ for each $k\in[K]$ and $\lambda$ satisfies (10), one can verify that the global solutions of Problem (5) satisfy (7) with $\overline{\sigma}_{1}=\dots=\overline{\sigma}_{K}=1$ and $\sum_{k=1}^{K}r_{k}=m$. Since $r_{k}\leq m_{k}$ for each $k\in[K]$, we have $r_{k}=m/K$ for each $k\in[K]$. Therefore, Problem (4) and Problem (5) have the same global solution set.
(ii) Suppose that $m\geq d$. Using this and (61), we have $\hat{\bm{Z}}_{k}=m\bm{U}_{k}\bm{V}_{k}^{T}/d$ and $\hat{r}_{k}=d/K$ for each $k\in[K]$. Moreover, according to Theorem 1, if $m_{k}=m/K$ for each $k\in[K]$ and $\lambda$ satisfies (9), one can verify that the global solutions of Problem (5) satisfy (7) with $\overline{\sigma}_{1}=\dots=\overline{\sigma}_{K}=m/d$ and $\sum_{k=1}^{K}r_{k}=d$. Therefore, the global solution set of Problem (4) is a subset of that of Problem (5).
@@
$\min_{\bm{x}\in\mathbb{R}^{r}}\sum_{i=1}^{r}-\log\left(1+\alpha x_{i}\right)+\sum_{i=1}^{r}\frac{1}{K}\log(1+K\alpha x_{i})\quad\mathrm{s.t.}\ \sum_{i=1}^{r}x_{i}=\frac{m}{K},\ x_{i}\geq 0,\ \forall i\in[r].$ (66)
@@
$f^{\prime}(x)=-\frac{\alpha}{1+\alpha x}+\frac{\alpha}{1+K\alpha x}<0,\quad f^{\prime\prime}(x)=\frac{\alpha^{2}}{(1+\alpha x)^{2}}-\frac{K\alpha^{2}}{(1+K\alpha x)^{2}}.$ (69)
@@
$\overline{x}=\frac{\eta+\sqrt{\Delta}}{2K\lambda\alpha},\quad\underline{x}=\frac{\eta-\sqrt{\Delta}}{2K\lambda\alpha}.$ (76)
@@
Case 2. Suppose that there exists $i\neq j\in\mathcal{S}$ such that $x_{i}\neq x_{j}$. This, together with $x_{i}\in\{\overline{x},\underline{x}\}$, implies $\overline{x}>\underline{x}$. According to (69), we have $f^{\prime\prime}(x)=0$ at $\hat{x}=1/(\alpha\sqrt{K})$. Then, we obtain that $f^{\prime}(x)$ is strictly decreasing in $[0,\hat{x}]$ and strictly increasing in $[\hat{x},\infty]$. Then, one can further verify that $\underline{x}<\hat{x}<\overline{x}$. This, together with (69), implies
@@
$f^{\prime\prime}(x_{r-1}^{\star})<0,\quad f^{\prime\prime}(x_{r}^{\star})<0.$ (79)
Using the second-order necessary condition for constraint optimization problems (see, e.g., [35, Theorem 12.5]) and $x_{i}^{\star}\geq 0$ for all $i\in[r]$, we obtain
@@
where the inequality is because $\overline{x}\leq{m}/\left({(s-1)K}\right)$ and $f(x)$ is strictly dereasing in $[0,\infty)$. For all $x\in\left[{m/(sK)},{m}/\left({(s-1)K}\right)\right]$, we compute
@@
$0\displaystyle 1+\frac{\log K}{K-1}-\log(1+\alpha)=\log\left(\frac{{K}^{\frac{1}{K-1}}\exp(1)}{1+\alpha}\right)<$ 0 . (85)
@@
This, together with Lemma 7, implies that when the optimal solution takes the form of (77) with $s=r$, Problem (66) achieves its global minimum. Then, we complete the proof. ∎
@@
Consider the setting in Lemma 6 and the following function
@@
D.3 Proof of and Theorem 2
Proof of Theorem 2.
According to (i) of Proposition 3, if $\bm{Z}$ is a critical point but not a local maximizer, we have $\bm{Z}\in\mathcal{Z}_{2}\cup\mathcal{Z}^{c}$. This, together with (ii) of Proposition 3 and Proposition 4, yields that $\bm{Z}$ is a strict saddle point. ∎
@@ Appendix E Additional Experimental Setups and Results
In this section, we provide additional implementation details and experimental results under different parameter settings for Sections 5.1 and 5.2 in Sections E.1 and E.2, respectively.
E.1 Implementation Details and Additional Results in Section 5.1
@@
In addition to the results presented in Section 5.1, we perform additional experiments under different settings as follows. To support our theorems, we visualize the heatmap of learned features and plot the number and magnitude of singular values in each class. Unless specified otherwise, we use the training setups introduced above in the following experiments.
@@
  • Experiment 1 on balanced data. In this experiment, we consider that the number of samples in each class is same and set the parameters in Problem (5) as follows: the dimension of features $d=100$, the number of classes $K=4$, the number of samples in each class is $m_{1}=m_{2}=m_{3}=m_{4}=50$, the regularization parameter $\lambda=0.1$, and the quantization error $\epsilon=0.5$. We visualize the heatmap between pairwise features of $\bm{Z}$ obtained by GD in Figure 5(a) and the number and magnitude of singular values in each class in Figure 6.
@@
  • Experiment 2 on data with more classes. In this experiment, we set the parameters in Problem (5) as follows: the dimension of features $d=100$, the number of classes $K=8$, the number of samples in each class is $m_{1}=50,m_{2}=40,m_{3}=30,m_{4}=20,m_{5}=30,m_{6}=40,m_{7}=40,m_{8}=50$, the regularization parameter $\lambda=0.1$, and the quantization error $\epsilon=0.5$. We visualize the heatmap between pairwise features of $\bm{Z}$ obtained by GD in Figure 5(b) and the number and magnitude of singular values in each class in Figure 7.
@@
  • Experiment 3 on data where the dimension $d$ is larger than the number of samples $m$. In this experiment, we set the parameters in Problem (5) as follows: the dimension of features $d=300$, the number of classes $K=4$, the number of samples in each class is $m_{1}=50,m_{2}=50,m_{3}=40,m_{4}=60$, the regularization parameter $\lambda=0.01$, and the quantization error $\epsilon=5$. Note that in this experiment, we set the learning rate as $1$. We visualize the heatmap between pairwise features of $\bm{Z}$ obtained by GD in Figure 5(c) and the number and magnitude of singular values in each class in Figure 8.
@@
According to the results in Figures 5(a), 5(b), 6, and 7 of Experiments 1 and 2, we observe that when the number of samples is larger than its dimension, i.e., $m\geq d$, the learned features via the MCR2 principle are within-class compressible and between-class discriminative in both balanced and unbalanced data sets. Moreover, the dimension of the space spanned by these features is maximized such that $\sum_{k=1}^{K}r_{k}=\min\{m,d\}$. This directly supports Theorem 1. By comparing the function value returned by GD and that computed by the closed-form in Theorem 1, we found that GD with random initialization will always converge to a global maximizer of Problem (5) when the data is balanced, while it will always converge to a local maximizer of Problem (5) when data is unbalanced. This directly supports Theorem 2.
@@
According to the results in Figures 5(c) and 8 of Experiment 3, we observe that when the number of samples is smaller than its dimension, i.e., $m\leq d$, the learned features via the MCR2 principle are orthogonal to each other and the dimension of each subspace is equal to the number of samples, i.e., $r_{k}=m_{k}$ for each $k\in[K]$. This exactly supports Theorem 1. Indeed, when $d\geq m$ and $r_{k}=m_{k}$ for each $k\in[K]$, it follows from Theorem 1 that $\bm{V}_{k}=\bm{I}$ and thus $\bm{Z}_{k}=\overline{\sigma}_{k}\bm{U}_{k}$ for each $k\in[K]$ for each local maximizer. Therefore, this also supports Theorem 2 as GD with random initialization converges to a local maximizer.
@@
Figure 9: Heatmap of cosine similarity among learned features by training deep networks on MNIST and CIFAR-10. We train network parameters by optimizing the regularized MCR2 objective (5) on $m$ samples split equally among $K$ classes of MNIST and CIFAR-10. In the figure, the darker pixels represent higher cosine similarity between features. In particular, when the $(i,j)$-th pixel is close to 00 (very light blue), the features $i$ and $j$ are approximately orthogonal.
E.2 Implementation Details and Additional Results in Section 5.2
@@
In the experiments on MNIST, we employ a 4-layer multilayer perception (MLP) network with ReLU activation as the feature mapping with the intermediate dimension 2048 and output dimension 32. In particular, each layer of MLP networks consists of a linear layer and layer norm layer followed by ReLU activation in the implementation. We train the network parameters via Adam by optimizing the MCR2 function. For the Adam settings, we use a momentum of 0.9, a full-batch size, and a dynamically adaptive learning rate initialized with $5\times 10^{-3}$, modulated by a CosineAnnealing learning rate scheduler [28]. We terminate the algorithm when it reaches 3000 epochs.
@@
Besides the numerical results in Table 1, we also plot the heatmap of the cosine similarity between pairwise columns of the features in $\bm{Z}$ obtained by training deep networks in Figure 9. We observe that the features from different classes are nearly orthogonal to each other, while those from the same classes are highly correlated. This supports our results in Theorem 1.
@@
In the experiments on CIFAR10, we employ a 2-layer multilayer perceptron (MLP) network with ReLU activation as the feature mapping with the intermediate dimension 4096 and output dimension 128. In particular, each layer of MLP networks consists of a linear layer and layer norm layer followed by ReLU activation in the implementation. We train the network parameters via Adam by optimizing the MCR2 function. For the Adam settings, we use a momentum of 0.9, a full-batch size, and a dynamically adaptive learning rate initialized with $5\times 10^{-3}$, modulated by a CosineAnnealing learning rate scheduler [28]. We terminate the algorithm when it reaches 4000 epochs.
@@
Besides the numerical results in Table 1, we also plot the heatmap of the cosine similarity between pairwise columns of the features in $\bm{Z}$ obtained by training deep networks in Figure 9. We observe that the features from different classes are nearly orthogonal to each other, while those from the same classes are highly correlated. This supports our results in Theorem 1.

### v2

@@ C.1 Proof of ˜3
where the second equality follows from $m_{k}\alpha_{k}/m=\alpha$ according to (11). This implies $\bm{0}$ is a local maximizer of $f_{k}(\bm{Z}_{k})$. Suppose to the contrary that $r_{k}>0$ for all $k\in[K]$. For each $\bm{Z}\in\cal Z$, using Lemma˜1 with $\bm{Z}_{k}^{T}\bm{Z}_{l}=\bm{0}$ for all $k\neq l$, (14), and (31), we have
@@
C.2 Proof of ˜4
Proof of ˜4.
@@ C.2 Proof of ˜4
According to Lemma˜5, we can see that $\bm{Z}$ is a critical point with curvature if and only if $\bm{Z}\bm{O}$ is a critical point with the same curvature for each $\bm{O}=\mathrm{BlkDiag}\left(\bm{O}_{1},\dots,\bm{O}_{K}\right)$ with $\bm{O}_{k}\in\mathcal{O}^{m_{k}}$ for all $k\in[K]$. According to the SVD of $\bm{Z}_{k}$ in (33), we can take $\bm{O}_{k}=\bm{Q}_{k}$ for each $k\in[K]$. Therefore, it suffices to study $\bm{Z}_{k}=\bm{P}_{k}\bm{\Sigma}_{k}$ for each $k\in[K]$. Substituting this into (24) in Definition˜1 gives
@@
This, together with (2), Lemma˜4, and $\bm{Z}^{(i)^{T}}\bm{Z}^{(j)}=\bm{0}$, yields that
@@
$\qquad-\alpha\alpha_{k}\left(1-\frac{\alpha_{k}}{\alpha_{k}\beta_{i}^{2}+1}\|\bm{Z}_{k}^{(i)^{T}}\bm{u}\|^{2}\right)\frac{\beta_{i}^{2}\|\bm{q}_{k}\|^{2}}{\alpha_{k}\beta_{i}^{2}+1},$
@@
$\qquad\sum_{k=1}^{K}\nabla^{2}R_{c}\left(\bm{Z}_{k}^{(i)}\right)\left[\bm{D}_{k}^{(i)},\bm{D}_{k}^{(i)}\right]$
@@
$\qquad-\frac{\alpha\alpha_{k}}{(\alpha_{k}\beta_{i}^{2}+1)^{2}}\sum_{k=1}^{K}\left(\bm{u}^{T}\bm{Z}_{k}^{(i)}\bm{q}_{k}\right)^{2}+\frac{\alpha\alpha_{k}^{2}\beta_{i}^{2}}{\left(\alpha_{k}\beta_{i}^{2}+1\right)^{2}}\sum_{k=1}^{K}\|\bm{q}_{k}\|^{2}\|\bm{Z}_{k}^{(i)^{T}}\bm{u}\|^{2}$
@@
Appendix D Proofs in Section˜3
D.1 Proof of Theorem˜1
@@
(i) Suppose that each block of $\bm{Z}$ satisfying (7), (a), (b), and (c). It directly follows from (i) in ˜3 that $\bm{Z}$ is a local maximizer. Conversely, suppose that $\bm{Z}$ is a local maximizer. According to (17), (18), (19), (ii) in ˜3 and ˜4, if $\bm{Z}\in\mathcal{Z}^{c}\cup\mathcal{Z}_{2}$, then $\bm{Z}$ is a strict saddle point. This, together with $\mathcal{X}=\mathcal{Z}^{c}\cup\mathcal{Z}_{1}\cup\mathcal{Z}_{2}$ and the fact that $\bm{Z}$ is a local maximizer, implies that $\bm{Z}\in\mathcal{Z}_{1}$. Using this, (19), and ˜2 yields that $\bm{Z}$ satisfying (7), (a), (b), and (c).
(ii) According to (i) in Theorem 1, suppose that the $k$-th block of a local maximizer $\bm{Z}$ admits the decomposition in (7) satisfying (a), (b), and (c) for all $k\in[K]$. This, together with (C.1) in the proof of ˜3, yields that
@@
D.2 Proof of ˜1
To prove ˜1, we first need to characterize the global optimal solution set of Problem (4).
@@ D.2 Proof of ˜1
According to Lemma˜4, we have
@@
$\max_{\sigma_{k,1},\dots,\sigma_{k,r_{k}}}\sum_{i=1}^{r_{k}}\log\left(1+\alpha\sigma_{k,i}^{2}\right)-\sum_{i=1}^{r_{k}}\frac{m_{k}}{m}\log\left(1+\alpha_{k}\sigma_{k,i}^{2}\right)\hskip 18.49988pt\mathrm{s.t.}\ \sum_{i=1}^{r_{k}}\sigma_{k,i}^{2}=m_{k}.$
@@
$\max_{x_{1},\dots,x_{r_{k}}}g(\bm{x})$ $:=\sum_{i=1}^{r_{k}}\log\left(1+\alpha x_{i}\right)-\sum_{i=1}^{r_{k}}\frac{1}{K}\log\left(1+\alpha_{k}x_{i}\right)\hskip 18.49988pt$
@@
This, together with Lemma˜6 and (11), yields that the optimal solution for each $k\in[K]$ is
@@
(i) Suppose that $m\leq d$. Then, we have $r_{k}\leq m/K$ for each $k\in[K]$ and $\sum_{k=1}^{K}r_{k}\leq m$. This, together with (65) and Lemma˜7, implies that $r_{k}=m/K$, and thus $x_{i}^{*}=1$ for all $i\in[r_{k}]$ and $k\in[k]$.
@@
Based on the above proposition, we are ready to prove ˜1.
Proof of ˜1.
Let $\hat{\bm{Z}}=[\bm{Z}_{1},\dots,\bm{Z}_{k}]$ denote the optimal solution of Problem (4). According to Proposition˜1, it suffices to study the following two cases.
(i) Suppose that $m<d$. Using this and (61), we have $\hat{\bm{Z}}_{k}=\bm{U}_{k}\bm{V}_{k}^{T}$ and $\hat{r}_{k}=m/K$ for each $k\in[K]$. Moreover, according to Theorem˜1, if $m_{k}=m/K$ for each $k\in[K]$ and $\lambda$ satisfies (10), one can verify that the global solutions of Problem (5) satisfy (7) with $\overline{\sigma}_{1}=\dots=\overline{\sigma}_{K}=1$ and $\sum_{k=1}^{K}r_{k}=m$. Since $r_{k}\leq m_{k}$ for each $k\in[K]$, we have $r_{k}=m/K$ for each $k\in[K]$. Therefore, Problem (4) and Problem (5) have the same global solution set.
(ii) Suppose that $m\geq d$. Using this and (61), we have $\hat{\bm{Z}}_{k}=m\bm{U}_{k}\bm{V}_{k}^{T}/d$ and $\hat{r}_{k}=d/K$ for each $k\in[K]$. Moreover, according to Theorem˜1, if $m_{k}=m/K$ for each $k\in[K]$ and $\lambda$ satisfies (9), one can verify that the global solutions of Problem (5) satisfy (7) with $\overline{\sigma}_{1}=\dots=\overline{\sigma}_{K}=m/d$ and $\sum_{k=1}^{K}r_{k}=d$. Therefore, the global solution set of Problem (4) is a subset of that of Problem (5).
@@
$\min_{\bm{x}\in\mathbb{R}^{r}}\sum_{i=1}^{r}-\log\left(1+\alpha x_{i}\right)+\sum_{i=1}^{r}\frac{1}{K}\log(1+K\alpha x_{i})\qquad\mathrm{s.t.}\ \sum_{i=1}^{r}x_{i}=\frac{m}{K},\ x_{i}\geq 0,\ \forall i\in[r].$ (66)
@@
$f^{\prime}(x)=-\frac{\alpha}{1+\alpha x}+\frac{\alpha}{1+K\alpha x}<0,\qquad f^{\prime\prime}(x)=\frac{\alpha^{2}}{(1+\alpha x)^{2}}-\frac{K\alpha^{2}}{(1+K\alpha x)^{2}}.$ (69)
@@
$\overline{x}=\frac{\eta+\sqrt{\Delta}}{2K\lambda\alpha},\qquad\underline{x}=\frac{\eta-\sqrt{\Delta}}{2K\lambda\alpha}.$ (76)
@@
Case 2. Suppose that there exists $i\neq j\in\mathcal{S}$ such that $x_{i}\neq x_{j}$. This, together with $x_{i}\in\{\overline{x},\underline{x}\}$, implies $\overline{x}>\underline{x}$. According to (69), we have $f^{\prime\prime}(x)=0$ at $\hat{x}=1/(\alpha\sqrt{K})$. Then, we obtain that $f^{\prime}(x)$ is strictly decreasing in $0,\hat{x}$] and strictly increasing in $\hat{x}$$\infty$]. Then, one can further verify that $\underline{x}<\hat{x}<\overline{x}$. This, together with (69), implies
@@
$f^{\prime\prime}(x_{r-1}^{\star})<0,\qquad f^{\prime\prime}(x_{r}^{\star})<0.$ (79)
Using the second-order necessary condition for constraint optimization problems (see, e.g., [nocedal1999numerical, Theorem 12.5]) and $x_{i}^{\star}\geq 0$ for all $i\in[r]$, we obtain
@@
where the inequality is because $\overline{x}\leq{m}/\left({(s-1)K}\right)$ and $f(x)$ is strictly dereasing in $0,\infty)$. For all $x\in\left[{m/(sK)},{m}/\left({(s-1)K}\right)\right]$, we compute
@@
$1+\frac{\log K}{K-1}-\log(1+\alpha)=\log\left(\frac{{K}^{\frac{1}{K-1}}\exp(1)}{1+\alpha}\right)<0.$ (85)
@@
This, together with Lemma˜7, implies that when the optimal solution takes the form of (77) with $s=r$, Problem (66) achieves its global minimum. Then, we complete the proof. ∎
@@
Consider the setting in Lemma˜6 and the following function
@@
D.3 Proof of and Theorem˜2
Proof of Theorem˜2.
According to (i) of ˜3, if $\bm{Z}$ is a critical point but not a local maximizer, we have $\bm{Z}\in\mathcal{Z}_{2}\cup\mathcal{Z}^{c}$. This, together with (ii) of ˜3 and ˜4, yields that $\bm{Z}$ is a strict saddle point. ∎
@@ Appendix E Additional Experimental Setups and Results
In this section, we provide additional implementation details and experimental results under different parameter settings for Sections˜5.1 and 5.2 in Sections˜E.1 and E.2, respectively.
E.1 Implementation Details and Additional Results in Section˜5.1
@@
In addition to the results presented in Section˜5.1, we perform additional experiments under different settings as follows. To support our theorems, we visualize the heatmap of learned features and plot the number and magnitude of singular values in each class. Unless specified otherwise, we use the training setups introduced above in the following experiments.
@@
  • Experiment 1 on balanced data. In this experiment, we consider that the number of samples in each class is same and set the parameters in Problem (5) as follows: the dimension of features $d=100$, the number of classes $K=4$, the number of samples in each class is $m_{1}=m_{2}=m_{3}=m_{4}=50$, the regularization parameter $\lambda=0.1$, and the quantization error $\epsilon=0.5$. We visualize the heatmap between pairwise features of $\bm{Z}$ obtained by GD in Figure˜5(a) and the number and magnitude of singular values in each class in Figure˜6.
@@
  • Experiment 2 on data with more classes. In this experiment, we set the parameters in Problem (5) as follows: the dimension of features $d=100$, the number of classes $K=8$, the number of samples in each class is $m_{1}=50,m_{2}=40,m_{3}=30,m_{4}=20,m_{5}=30,m_{6}=40,m_{7}=40,m_{8}=50$, the regularization parameter $\lambda=0.1$, and the quantization error $\epsilon=0.5$. We visualize the heatmap between pairwise features of $\bm{Z}$ obtained by GD in Figure˜5(b) and the number and magnitude of singular values in each class in Figure˜7.
@@
  • Experiment 3 on data where the dimension $d$ is larger than the number of samples $m$. In this experiment, we set the parameters in Problem (5) as follows: the dimension of features $d=300$, the number of classes $K=4$, the number of samples in each class is $m_{1}=50,m_{2}=50,m_{3}=40,m_{4}=60$, the regularization parameter $\lambda=0.01$, and the quantization error $\epsilon=5$. Note that in this experiment, we set the learning rate as $1$. We visualize the heatmap between pairwise features of $\bm{Z}$ obtained by GD in Figure˜5(c) and the number and magnitude of singular values in each class in Figure˜8.
@@
According to the results in Figures 5(a), 5(b), 6, and 7 of Experiments 1 and 2, we observe that when the number of samples is larger than its dimension, i.e., $m\geq d$, the learned features via the MCR2 principle are within-class compressible and between-class discriminative in both balanced and unbalanced data sets. Moreover, the dimension of the space spanned by these features is maximized such that $\sum_{k=1}^{K}r_{k}=\min\{m,d\}$. This directly supports Theorem˜1. By comparing the function value returned by GD and that computed by the closed-form in Theorem˜1, we found that GD with random initialization will always converge to a global maximizer of Problem (5) when the data is balanced, while it will always converge to a local maximizer of Problem (5) when data is unbalanced. This directly supports Theorem˜2.
@@
According to the results in Figures 5(c) and 8 of Experiment 3, we observe that when the number of samples is smaller than its dimension, i.e., $m\leq d$, the learned features via the MCR2 principle are orthogonal to each other and the dimension of each subspace is equal to the number of samples, i.e., $r_{k}=m_{k}$ for each $k\in[K]$. This exactly supports Theorem˜1. Indeed, when $d\geq m$ and $r_{k}=m_{k}$ for each $k\in[K]$, it follows from Theorem˜1 that $\bm{V}_{k}=\bm{I}$ and thus $\bm{Z}_{k}=\overline{\sigma}_{k}\bm{U}_{k}$ for each $k\in[K]$ for each local maximizer. Therefore, this also supports Theorem˜2 as GD with random initialization converges to a local maximizer.
@@
CIFAR: $m=2500,K=10$
Figure 9: Heatmap of cosine similarity among learned features by training deep networks on MNIST and CIFAR-10. We train network parameters by optimizing the regularized MCR2 objective (5) on $m$ samples split equally among $K$ classes of MNIST and CIFAR-10. In the figure, the darker pixels represent higher cosine similarity between features. In particular, when the $(i,j)$-th pixel is close to 0 (very light blue), the features $i$ and $j$ are approximately orthogonal.
E.2 Implementation Details and Additional Results in Section˜5.2
@@
In the experiments on MNIST, we employ a 4-layer multilayer perception (MLP) network with ReLU activation as the feature mapping with the intermediate dimension 2048 and output dimension 32. In particular, each layer of MLP networks consists of a linear layer and layer norm layer followed by ReLU activation in the implementation. We train the network parameters via Adam by optimizing the MCR2 function. For the Adam settings, we use a momentum of 0.9, a full-batch size, and a dynamically adaptive learning rate initialized with $5\times 10^{-3}$, modulated by a CosineAnnealing learning rate scheduler [loshchilov2016sgdr]. We terminate the algorithm when it reaches 3000 epochs.
@@
Besides the numerical results in Table˜1, we also plot the heatmap of the cosine similarity between pairwise columns of the features in $\bm{Z}$ obtained by training deep networks in Figure˜9. We observe that the features from different classes are nearly orthogonal to each other, while those from the same classes are highly correlated. This supports our results in Theorem˜1.
@@
In the experiments on CIFAR10, we employ a 2-layer multilayer perceptron (MLP) network with ReLU activation as the feature mapping with the intermediate dimension 4096 and output dimension 128. In particular, each layer of MLP networks consists of a linear layer and layer norm layer followed by ReLU activation in the implementation. We train the network parameters via Adam by optimizing the MCR2 function. For the Adam settings, we use a momentum of 0.9, a full-batch size, and a dynamically adaptive learning rate initialized with $5\times 10^{-3}$, modulated by a CosineAnnealing learning rate scheduler [loshchilov2016sgdr]. We terminate the algorithm when it reaches 4000 epochs.
@@
Besides the numerical results in Table˜1, we also plot the heatmap of the cosine similarity between pairwise columns of the features in $\bm{Z}$ obtained by training deep networks in Figure˜9. We observe that the features from different classes are nearly orthogonal to each other, while those from the same classes are highly correlated. This supports our results in Theorem˜1.